OpenStudy (anonymous):

Help please? 7y + 5 - 3y + 1 = 2y + 2

OpenStudy (anonymous):

7y + 5 - 3y + 1 = 2y + 2 combine similar terms first

OpenStudy (anonymous):

7y-3y-2y=2-5-1 then solve now ^_^ 2y=-4 (2y=-4)/2 y=-2